OpenStudy (anonymous):

Which of the following statements is false? A)The product of two rational numbers is always irrational. B)The sum of a rational number and an irrational number is always irrational. C)The product of a nonzero rational number and an irrational number is always irrational. D)The product of two irrational numbers is either rational or irrational.

OpenStudy (anonymous):

@Astrophysics

OpenStudy (anonymous):

@SurpriseBreeze

OpenStudy (unklerhaukus):

you were meant to find the option that is FALSE

OpenStudy (unklerhaukus):

an example of The product of a non-zero rational, and a irrational number, begin equal to an irrational " 2 x √3 = 2√3 C is not FALSE

OpenStudy (mrnood):

\[\frac{ 2 }{ 1 }\times \frac{ 3 }{ 1 } =??\] Product of 2 rational numbers...